OpenStudy (anonymous):

Find dy/dx: y^2-xy=3

OpenStudy (anonymous):

differentiating wrt x we get 2ydy/dx - y - xdy/dx = 0 = dy/dx = y/(2y + x)

OpenStudy (ash2326):

differentiate with respect to x 2ydy/dx-xdy/dx-y=0 dy/dx(2y-x)-y=0 dy/dx=y/(2y-x)